OpenStudy (andijo76):

mixed nuts sell pistachio kernals for $10.00 per pound and almonds for $8.00 per pound how much of each type should be used to make a 50lb mixture that sells for 8.80 per pound

OpenStudy (ranga):

Let the mixture have x pounds of pistachio and y pounds of almonds. x + y = 50 --- (1) 10x + 8y = 8.80 * 50 or 10x + 8y = 440 or x + 0.8y = 44 --- (2) Solve for x and y. Subtract (2) from (1) and you can solve for y. Then put it in (1) and solve for x.

OpenStudy (ranga):

If any step needs clarification please let me know.

OpenStudy (andijo76):

i dont know if i am correct does it come out to 84 bags

OpenStudy (ranga):

x + y = 50 --- (1) x + 0.8y = 44 --- (2) subtract 0.2y = 6 y = 6/0.2 = 30 put y = 30 in (1) x + 30 = 50 x = 20 20 lbs of pistachios should be mixed with 30 lbs of almonds.

OpenStudy (andijo76):

thank you

OpenStudy (ranga):

you are welcome.